# Model-Based Learning for DOA Estimation with One-Bit Single-Snapshot Sparse Arrays

### v1

@@ IV-A1 The SBRI-Net Framework
where $F_{conv}^{k}$ represents the CNN structure in the $k$-th layer, and $\mathbf{W}_{conv}^{k}$ and $\mathbf{b}_{conv}^{k}$ denote its corresponding weights and biases, respectively. Assuming that the CNN structure consists of $J$ convolutional layers, the total number of convolutional layers is $KJ$.
@@
where $\mathbf{W}_{FC}^{k}$ and $\mathbf{b}_{FC}^{k}$ denote the weights and bias, respectively, for the $k$-th fully connected layer. Instead of updating the grid gaps as described in line 6 of Algorithm 2, we utilize fully connected layers to produce the grid gap estimates for the next layer.
@@ IV-B Data Generation and Labeling
We generate training datasets for on-grid and off-grid DOA estimation tasks, respectively using two SLA configurations with half-wavelength element spacing: an 18-element array at $\frac{\lambda}{2}[0,1,2,3,4,7,8,9,10,11,12,13,14,15,16,17,18,19]$ and a 10-element array at $\frac{\lambda}{2}[0,3,4,5,6,7,11,16,18,19]$. The number of signals is set to 2, and the array field of view (FOV) is defined as $[-60^{\circ},60^{\circ}]$. For on-grid training data generation, the target DOAs are randomly sampled as integer degrees within the FOV. For off-grid training data generation, the angular space is discretized into uniform grids with a fixed grid size of $2^{\circ}$. The off-grid gaps corresponding to the signals are randomly generated following a uniform distribution $U\left(-1^{\circ},1^{\circ}\right)$. Reflection coefficients of the two signals are generated as random complex numbers, with their real and imaginary parts uniformly distributed in $U\left(0.5,1\right)$. Denoting the ground truth of the $n$-th DOA as set ${\mathbb{G}}_{n}$, the signals are labeled according to
@@
We randomly generate $1,000,000$ samples across input SNR levels ranging between 00 dB and $30$ dB in $5$ dB increments. $90\%$ of the dataset is used for training and the remaining $10\%$ is used for validation.
@@ V Numerical Results
To analyze the performance of the proposed on-grid and off-grid algorithm frameworks, we use root mean square error (RMSE) and hit rate as the evaluation metrics. The RMSE is computed as $\mathrm{RMSE}=\sqrt{\frac{1}{N_{s}{K}}\sum_{t=1}^{N_{s}}||\hat{\bm{\theta}}_{t}-\bm{\theta}^{\star}||_{2}^{2}}$, where $\hat{\bm{\theta}}_{t}$ represents the estimated DOA vector in the $t$-th test round, $K$ is the number of signals, and $N_{s}$ is the number of successful tests. The hit rate is defined as $\mathrm{Hit\ rate}={N_{s}}/{N_{t}}$, where $N_{t}$ is the total number of simulation trails. A DOA estimate is considered successful if the absolute errors of all estimated angles remain within a threshold of $2^{\circ}$; otherwise, it is regarded as a failure.
@@ V-A On-Grid DOA Estimation
We conducted $1,000$ simulation trials using an 18-element SLA to compare the average RMSE and hit rate of the proposed algorithm against the 1-bit SLIM method [26]. Figure 6(a) presents the estimation RMSEs of the proposed 1-bit SBRI and 1-bit SBRI-X, alongside 1-bit SLIM, across different input SNR levels. Figure 6 (b) illustrates the estimation hit rate for these methods. The test scenario consists of two fixed targets located at $\left[-30^{\circ},30^{\circ}\right]$. The proposed 1-bit SBRI-X achieves the lowest RMSE across all SNR levels, outperforming both 1-bit SLIM and 1-bit SBRI. Additionally, it demonstrates the highest estimation accuracy and highest hit rate in low input SNR conditions between 0 dB and 15 dB. These results highlight the performance gains of 1-bit SBRI-X over 1-bit SBRI and validate that the chosen sigmoid likelihood function contributes significantly to performance gains, particularly in low SNR conditions.
(a) RMSE vs. input SNR
(b) Hit rate vs. input SNR
Figure 6: Performance comparison for on-grid DOA estimation using $18$-element SLA.
@@
We further investigate the impact of parameter selection on the performance of the proposed 1-bit SBRI-X algorithm. As shown in Figure 1, parameter $\mathrm{b}$ controls the slopes of the likelihood functions. Figure 7 presents the estimation curves of 1-bit SBRI-X under different $\mathrm{b}$ settings. It can be observed that reducing $\mathrm{b}$ results in lower RMSEs across various SNR levels and a higher hit rate, indicating that smaller slopes of the likelihood function enhance the performance. It is also clear from Figure 7(a) that the RMSE initially decreases and then increases as the input SNR increases. This trend can be attributed to the influence of noise and quantization error at different SNR levels. In low SNR conditions, noise is the dominant factor affecting estimation accuracy, whereas in high SNR conditions, the impact of noise diminishes, and quantization error becomes the dominant factor. As shown in the results of Figure 8, the estimation errors decrease as the SNR increases and eventually converge. Additionally, Figure 8 presents the convergence curves for different choices of $\mathrm{b}$. It can be observed that smaller values of $\mathrm{b}$ (e.g., $\mathrm{b}=0.1$ or $\mathrm{b}=0.5$) result in a faster convergence rate, reaching smaller approximation error within approximately $10$ iterations.
(a) RMSE vs. input SNR
(b) Hit rate vs. input SNR
Figure 7: Effects of parameter selection on the performance of the proposed augmented algorithm framework for on-grid DOA estimation with an 18-element SLA.
@@
To further evaluate the performance of the proposed SBRI-Net and SBRI-X-Net (illustrated in Figures 2 and 3), we conducted $5,000$ random trials using an 18-element SLA. These trials focused on a scenario with two fixed on-grid targets located at $\left[-10^{\circ},20^{\circ}\right]$ under varying SNR levels. The results were compared against optimization-based estimators, including 1-bit SPICE [26], 1-bit LIKES [26], and 1-bit SLIM [26]. As shown in Figures 9(a) and 9(b), SBRI-X-Net achieves comparable RMSE to SBRI-X at SNRs from 00 dB to $15$ dB, and slightly lower RMSE at $20$ dB to $50$ dB. It also maintains a comparable hit rate across various SNR conditions, including those ranging from 35 dB to 50 dB, which were not part of the training scenarios. This demonstrates better overall performance and a certain level of generalization ability in high SNR conditions. A similar trend is observed for SBRI-Net, further demonstrating the effectiveness of the proposed model-based networks.
Figure 8: Convergence curves of 1-bit SBRI-X.
(a) RMSE vs. input SNR
(b) Hit rate vs. input SNR
@@
Figure 9: The performance of the proposed augmented algorithms and networks for on-grid DOA estimation with an 18-element SLA.
@@ V-B Off-Grid DOA Estimation
We first conducted $5,000$ simulation trials on two fixed off-grid targets at $\left[-10.28^{\circ},20.56^{\circ}\right]$, using the 18-element SLA described in Section IV-B. The proposed algorithms and networks were evaluated against several benchmarks, including the pure data-driven CNN-DNN-based method proposed in [51], and other algorithms such as OGIR [35], 1-bit SLIM, 1-bit LIKES and 1-bit SPICE [26]. As shown in Figure 10(a), the off-grid methods consistently outperform on-grid algorithms. Among off-grid approaches, SBRI-X-Net achieved the lowest RMSE, while SBRI-Net performed on par with OGIR. Furthermore, the proposed model-based SBRI-X-Net and SBRI-Net exhibit superior performance over the pure data-driven CNN-DNN approach. From Figure 10(b), the proposed model-based approach achieves the highest hit rate across various SNR conditions, confirming the superior estimation performance of the proposed methods.
(a) RMSE vs. input SNR
(b) Hit rate vs. input SNR
Figure 10: Performance comparison of SBRI-Net for off-grid DOA estimation using an $18$-element SLA.
(a) RMSE vs. input SNR
(b) Hit rate vs. input SNR
Figure 11: Performance comparison of SBRI-Net for off-grid DOA estimation using an $10$-element SLA.
In Figure 11, we further compare the performance of the proposed networks using the 10-element SLA described in Section IV-B. It can be observed that all methods experience performance degradation on the sparser 10-element SLA compared to their results with the 18-element SLA. However, the proposed model-based algorithms still outperform the other comparison algorithms, demonstrating their robustness for SLA configurations.
@@ VI Conclusions
The derivations of the updating steps in Equations (24) and (25) are as follows. First, we compute the first-order derivative of the objective function in Equation (39) with respect to $\hat{\mathbf{x}}$:

### v2

@@ IV-A1 The SBRI-Net Framework
where $F_{conv}^{k}$ represents the CNN structure in the $k$-th layer, and $\mathbf{W}_{conv}^{k}$ and $\mathbf{b}_{conv}^{k}$ denote its corresponding weights and biases, respectively. Assuming that the CNN structure consists of $J$ 1D convolutional layers, the total number of convolutional layers is $KJ$, in our experiments we set $J=4$.
@@
where $\mathbf{W}_{FC}^{k}$ and $\mathbf{b}_{FC}^{k}$ denote the weights and bias, respectively, for the $k$-th fully connected layer. Instead of updating the grid gaps as described in line 7 of Algorithm 2, we utilize fully connected layers to produce the grid gap estimates for the next layer.
@@ IV-B Data Generation and Labeling
We generate training datasets for on-grid and off-grid DOA estimation tasks, respectively using two SLA configurations with half-wavelength element spacing: an 18-element array at $\frac{\lambda}{2}[0,1,2,3,4,7,8,9,10,11,12,13,14,15,16,17,18,19]$ and a 10-element array at $\frac{\lambda}{2}[0,3,4,5,6,7,11,16,18,19]$. The number of signals is set to 2, and the array field of view (FOV) is defined as $60^{\circ},60^{\circ}]$. For on-grid training data generation, the target DOAs are randomly sampled as integer degrees within the FOV. For off-grid training data generation, the angular space is discretized into uniform grids with a fixed grid size of $2^{\circ}$. The off-grid gaps corresponding to the signals are randomly generated following a uniform distribution $U\left(-1^{\circ},1^{\circ}\right)$. Reflection coefficients of the two signals are generated as random complex numbers, with their real and imaginary parts uniformly distributed in $U\left(0.5,1\right)$. Denoting the ground truth of the $n$-th DOA as set ${\mathbb{G}}_{n}$, the signals are labeled according to
@@
We randomly generate $1,000,000$ samples across input SNR levels ranging between 0 dB and $30$ dB in $5$ dB increments. $90\%$ of the dataset is used for training and the remaining $10\%$ is used for validation.
@@ IV-D Network Training
IV-E Complexity Analysis and Comparison of Different One-Bit DOA Estimation Algorithms
The computational complexity of the proposed 1-bit on-grid and off-grid algorithms can be summarized as follows. For SBRI, initialization requires approximately $O(NM)$ floating-point operations (FLOPs). Each iteration is dominated by a matrix inversion of size $N$, leading to a per-iteration complexity of about $\frac{2}{3}N^{3}+O(MN^{2}+NM)$ FLOPs. Therefore, the overall complexity across $N_{\rm iter}$ iterations is $O(N_{\rm iter}N^{3})$. Off-grid SBRI follows a similar pattern, as the additional grid-gap update step does not change the asymptotic cubic scaling.
For SBRI-X, both the on-grid and off-grid variants maintain the same initialization cost as SBRI, while each iteration incurs similar computational demands due to matrix inversion and auxiliary variable updates. Consequently, the overall complexity for SBRI-X also scales as $O(N_{\rm iter}N^{3})$. Thus, all classical and augmented iterative algorithms exhibit cubic scaling with respect to the grid size $N$.
For the inference network SBRI-Net, the computational complexity for 1-bit on-grid DOA estimation is approximately $8NM$ FLOPs for initialization. Each of the $K$ unrolled layers then involves an $8NM$ auxiliary-vector update and 13,088N13{,}088N FLOPs for the convolution block $F_{conv}^{k}$ in (61). The parameters associated with the input channels, $C_{in}$, output channels, $C_{out}$, and channel length, $K_{in}$, are given as $(C_{in},C_{out},K_{in})=\{(1,21,21),(21,11,11),(11,5,5),(5,1,3)\}$. This yields a total complexity of K(8NM+13,088N)K(8NM+13{,}088N), which is $O(KN)$ for moderate $M$. For off-grid SBRI-Net, initialization requires $16NM$ FLOPs, and each layer incurs $16NM$ for the auxiliary update, 13,088N13{,}088N for convolution, and $12N^{2}$ for three fully connected layers, giving $K(16NM+13{,}088N+12N^{2})$ FLOPs in total. Thus, the asymptotic complexity is $O(KN)$ for on-grid, and $O(KN^{2})$ for off-grid estimation.
For SBRI-X-Net, the complexity follows the same structure: on-grid estimation requires K(12NM+13,088N)K(12NM+13{,}088N) FLOPs, scaling linearly with $N$, while off-grid estimation needs $K(20NM+13{,}088N+12N^{2})$ FLOPs per $K$ layers, growing quadratically with $N$ due to the fully connected blocks.
Table I summarizes the complexity of the proposed methods and reference methods such as 1-bit LIKES, 1-bit SPICE, 1-bit SLIM[28, 58].
Table I: Computational complexity of the compared DOA estimators ($M$ sensors, NN grid points, $N_{\mathrm{iter}}$ iterations, $K$ unrolled layers).
Algorithm Total cost
1-bit LIKES $\tfrac{5}{3}N_{\mathrm{iter}}M^{3}\;+\;4\,N_{\mathrm{iter}}M^{2}N$
1-bit SPICE $\tfrac{5}{3}N_{\mathrm{iter}}M^{3}\;+\;2\,N_{\mathrm{iter}}M^{2}N$
1-bit SLIM $\tfrac{5}{3}N_{\mathrm{iter}}M^{3}\;+\;4\,N_{\mathrm{iter}}M^{2}N$
1-bit SBRI (on-grid) $\tfrac{2}{3}N_{\mathrm{iter}}N^{3}\;+\;2\,N_{\mathrm{iter}}MN^{2}$
1-bit SBRI (off-grid) $N_{\mathrm{iter}}N^{3}\;+\;3\,N_{\mathrm{iter}}MN^{2}$
1-bit SBRI-Net (on-grid) K(8NM+ 13 088N)K\,(8\,N\,M\;+\;13\,088\,N)
1-bit SBRI-Net (off-grid) $K\,(16\,N\,M\;+\;13\,088\,N\;+\;12\,N^{2})$
Table II presents a comparison of the key features of different algorithms for 1-bit DOA estimation.
Table II: Key features comparison for different DOA estimation approaches.
Algorithms Snapshot SLA 1-Bit Off-Grid
1-bit SLIM Single
1-bit SPICE Single
1-bit LIKES Single
OGIR Single/Multiple
1-bit on-grid SBRI Single
1-bit off-grid SBRI Single
1-bit on-grid SBRI-X Single
1-bit off-grid SBRI-X Single
1-bit SBRI-Net Single/Multiple
1-bit SBRI-X-Net Single/Multiple
@@ V Numerical Results
[…]
@@ V-A On-Grid DOA Estimation
We conducted $1,000$ simulation trials using an 18-element SLA to compare the average RMSE and hit rate of the proposed algorithm against the 1-bit SLIM method [28]. Figure 11(a) (a) presents the estimation RMSEs of the proposed 1-bit SBRI and 1-bit SBRI-X, alongside 1-bit SLIM, across different input SNR levels. Figure 11(a) (b) illustrates the estimation hit rate for these methods. The test scenario consists of two fixed targets located at $\left[-30^{\circ},30^{\circ}\right]$. The proposed 1-bit SBRI-X achieves the lowest RMSE across all SNR levels, outperforming both 1-bit SLIM and 1-bit SBRI. Additionally, it demonstrates the highest estimation accuracy and highest hit rate in low input SNR conditions between 0 dB and 15 dB. These results highlight the performance gains of 1-bit SBRI-X over 1-bit SBRI and validate that the chosen sigmoid likelihood function contributes significantly to performance gains, particularly in low SNR conditions.
(a) Performance comparison for on-grid DOA estimation using $18$-element SLA: (a) RMSE versus input SNR; (b) Hit rate versus input SNR; (c) RMSE versus input SNR with different selection of $\mathrm{a}$ (fixed $\mathrm{b}$); (d) Hit rate versus input SNR with different selection of $\mathrm{a}$ (fixed $\mathrm{b}$); (e) RMSE versus input SNR with different selection of $\mathrm{b}$ (fixed $\mathrm{a}$); (f) Hit rate versus input SNR with different selection of $\mathrm{b}$ (fixed $\mathrm{a}$).
@@
We further investigate the impact of parameter selection on the performance of the proposed 1-bit SBRI-X algorithm. As shown in Figure 1, parameter $\mathrm{b}$ controls the slopes of the likelihood functions. Figure 11(a) (c-d) presents the estimation curves of 1-bit SBRI-X under different $\mathrm{b}$ settings. It can be observed that reducing $\mathrm{b}$ results in lower RMSEs across various SNR levels and a higher hit rate, indicating that smaller slopes of the likelihood function enhance the performance. Figure 11(a) (c) shows that the RMSE decreases at first and then rises as the input SNR increases. This trend can be attributed to the influence of noise and quantization error at different SNR levels. In low SNR conditions, noise is the dominant factor affecting estimation accuracy, whereas in high SNR conditions, the impact of noise diminishes, and quantization error becomes the dominant factor. This performance saturation phenomenon has also been reported in the literature, e.g., [18]. To conduct more ablation studies, we compared estimation performance under fixed $b=1.0$ but various $\mathrm{a}$. Figure 11(a) (e) shows that increasing $\mathrm{a}$ reduces RMSE at every SNR, but the gain saturates when $\mathrm{a}\geq 3$; beyond this point the curve flattens, demonstrating a “diminishing-returns” effect frequently reported in hyper-parameter studies. In low-SNR regimes (01010 dB), larger $\mathrm{a}$ gives a marked RMSE improvement, whereas at high SNR an overly larger $\mathrm{a}$ slightly degrades accuracy. Hit-rate curves in Figure 11(a) (f) mirror the same trend: increasing $a$ maximises detections. Taken together with the sensitivity to $b$ shown in Figure 11(a) (e-f), these results confirm that SBRI-X is responsive to $\left(\mathrm{a},\mathrm{b}\right)$. Together, these ablations confirm that both $\mathrm{a}$ and $\mathrm{b}$ shape SBRI-X performance.
We can further optimize the parameters $\mathrm{a}$ and $\mathrm{b}$ via two-dimensional (2D) grid search. As illustrated in Figure 11(b), the lowest RMSE is attained when the parameter pair $\left(\mathrm{a},\mathrm{b}\right)$ is close to $\left(0.5,1.7\right)$.
(b) Effects of 2D tuning of the sigmoid parameters $\mathrm{a}$ and $\mathrm{b}$ on RMSE and Hit rate for on-grid DOA estimation using the proposed SBRI-X (18-element SLA) under 10 dB input SNR: (a) RMSE versus 2D grid search with different $\mathrm{a}$ and $\mathrm{b}$; (b) Hit rate versus 2D grid search with different $\mathrm{a}$ and $\mathrm{b}$; (c) Convergence curves of 1-bit SBRI-X.
Additionally, Figure 11(b) (c) presents the convergence curves for different choices of $\mathrm{b}$. It can be observed that smaller values of $\mathrm{b}$ (e.g., $\mathrm{b}=0.1$ or $\mathrm{b}=0.5$) result in a faster convergence rate, reaching smaller approximation error within approximately $10$ iterations.
Note also that the choice of the regularization parameter $\gamma$ will affect the algorithm convergence speed. We compared convergence under different $\gamma$ settings and found that, as Figure 11(c) (a) shows, the adaptive rule in Equation (20) accelerates convergence compared with a fixed $\gamma$.
@@
To further evaluate the performance of the proposed SBRI-Net and SBRI-X-Net (illustrated in Figure 2 and Figure 3), we conducted $5,000$ random trials using an 18-element SLA. These trials focused on a scenario with two fixed on-grid targets located at $\left[-10^{\circ},20^{\circ}\right]$ under varying SNR levels. The results were compared against optimization-based estimators, including 1-bit SPICE, 1-bit LIKES, and 1-bit SLIM [28]. As shown in Figure 11(c) (b) and (c), SBRI-X-Net achieves comparable RMSE to SBRI-X at SNRs from 0 dB to $15$ dB, and slightly lower RMSE at $20$ dB to $50$ dB. It also maintains a comparable hit rate across various SNR conditions, including those ranging from $35$ dB to $50$ dB, which were not part of the training scenarios. This demonstrates better overall performance and a certain level of generalization ability in high SNR conditions. A similar trend is observed for SBRI-Net, further demonstrating the effectiveness of the proposed model-based networks.
@@
(c) The performance of the proposed augmented algorithms and networks for on-grid DOA estimation with an 18-element SLA: (a) Convergence curves of 1-bit SBRI with different regularization settings; (b) RMSE versus input SNR; (c) Hit rate versus input SNR.
Unlike the original SBRI-X algorithm, which requires manually chosen $a$ and $b$, SBRI-X-Net learns these parameters during training. We initialize both $a$ and $b$ to 1.0. Because SBRI-X is not strictly an unrolled network, letting both parameters float simultaneously led to numerical instabilities (NaN propagation). To avoid this, we fix $a$ and allow only $b$ to be trainable. Figure 11 plots $b$ (and fixed $a$) over training epochs. We see that $b$ steadily decreases, mirroring the behavior observed in our manual grid searches (Figure 11(a) (c-d)), where smaller $b$ consistently yielded better estimation performance.
Figure 11: Evolution of the learned sigmoid parameters $(\mathrm{a},\mathrm{b})$ across training epochs for SBRI-X-Net.
@@ V-B Off-Grid DOA Estimation
We first conducted $5,000$ simulation trials on two fixed off-grid targets at $\left[-10.28^{\circ},20.56^{\circ}\right]$, using the 18-element SLA described in Section IV-B. The proposed algorithms and networks were evaluated against several benchmarks, including the pure data-driven CNN-DNN-based method proposed in [59], and other algorithms such as OGIR [37], 1-bit SLIM, 1-bit LIKES and 1-bit SPICE [28]. As shown in Figure 17(a) (a), the off-grid methods consistently outperform on-grid algorithms. Among off-grid approaches, SBRI-X-Net achieved the lowest RMSE, while SBRI-Net performed on par with OGIR. Furthermore, the proposed model-based SBRI-X-Net and SBRI-Net exhibit superior performance over the pure data-driven CNN-DNN approach. From Figure 17(a) (b), the proposed model-based approach achieves the highest hit rate across various SNR conditions, confirming the superior estimation performance of the proposed methods.
(a) Performance comparison of SBRI-Net for off-grid DOA estimation: (a) RMSE versus input SNR for $18$-element SLA with two targets; (b) Hit rate versus input SNR for $18$-element SLA with two targets; (c) RMSE versus input SNR for an $10$-element SLA with two targets; (d) Hit rate versus input SNR for $10$-element SLA with two targets; (e) RMSE versus input SNR for $10$-element SLA with three targets; (f) Hit rate versus input SNR for $10$-element SLA with three targets.
In Figure 17(a) (c-d), we further compare the performance of the proposed networks using the 10-element SLA described in Section IV-B. It can be observed that all methods experience performance degradation on the sparser 10-element SLA compared to their results with the 18-element SLA. However, the proposed model-based algorithms still outperform the other comparison algorithms, demonstrating their robustness for SLA configurations.
To further assess robustness at the scenario containing more sources, we carried out an additional 5,0005{,}000 Monte-Carlo trials with three off–grid targets positioned at $(-10.28^{\circ},\,0.25^{\circ},\,20.56^{\circ})$. As summarized in Figure 17(a) (e-f), the trends remain consistent with the two–source experiment: SBRI-X-Net delivers the lowest RMSE across the entire SNR range, while SBRI-Net maintains clear gains over OGIR (especially under low input SNR) and the pure CNN-DNN baseline. Both networks also achieve the highest hit rate, demonstrating that the proposed framework scales gracefully to scenarios with more active sources.
@@ VI Conclusions
Appendix A Proof of Theorem 1
Proof.
For $\eta>0$, the coercivity of the regularizer ensures that the sequence $\{\hat{\mathbf{x}}^{k}\}_{k\geq 0}$ remains bounded. Hence, according to the Bolzano-Weierstrass theorem[60], there exists a convergent subsequence $\{\hat{\mathbf{x}}^{k_{\jmath}}\}$ such that
$\hat{\mathbf{x}}^{k_{\jmath}}\to\hat{\mathbf{x}}^{\dagger},\quad\text{as }\jmath\to\infty.$ (78)
By Lemma 2 (see Appendix C), we have
$\|\hat{\mathbf{x}}^{k+1}-\hat{\mathbf{x}}^{k}\|_{2}\to 0,$ (79)
which implies that the shifted subsequence $\{\hat{\mathbf{x}}^{k_{\jmath+1}}\}$ also converges to the limit $\hat{\mathbf{x}}^{\dagger}$. Let $\{\mathbf{v}^{k_{\jmath+1}}\}$ denote the corresponding subsequence of the $\mathbf{v}$-iterates. Since the sequence $\{(\hat{\mathbf{x}}^{k},\mathbf{v}^{k})\}$ is bounded and from Lemma 3 (see Supplementary Materials), the Lyapunov function is monotonically decreasing, we define the limit point:
$\mathbf{v}^{\dagger}:=\lim_{\jmath\to\infty}\mathbf{v}^{k_{\jmath}}.$ (80)
Letting $\jmath\to\infty$ in the first-order optimality condition associated with the $\hat{\mathbf{x}}-$update (Equation (18)), we obtain:
$\mathbf{z}+\mathcal{A}^{\mathrm{H}}(\mathcal{A}\hat{\mathbf{x}}^{\dagger}-\mathbf{v}^{\dagger})=\mathbf{0},$ (81)
where ${\mathbf{z}}=[z_{1},\cdots,z_{N}]^{T}$ is the gradient of the smoothed regularizer:
$z_{i}=\frac{\gamma\hat{x}_{i}^{\dagger}}{(\eta+|\hat{x}_{i}^{\dagger}|^{2})^{1-\frac{\alpha}{2}}},\ i=1,...,N,$ (82)
and $\hat{\mathbf{x}}^{\dagger}$ is the limit point of the surrogate objective in the $\hat{\mathbf{x}}-$update step. Furthermore, Lemma 2 in the Appendix C implies that the function $\mathcal{L}(\hat{\mathbf{x}}^{k},\gamma)$ is monotonically decreasing and bounded below, so it converges. In particular, we have
$\mathcal{L}(\hat{\mathbf{x}}^{\dagger},\gamma)\leq\mathcal{L}(\hat{\mathbf{x}}^{0},\gamma)\leq\|\hat{\mathbf{x}}^{0}\|_{\alpha}^{2}+N,$ (83)
where $\|\hat{\mathbf{x}}\|_{\alpha}^{2}:=\sum_{i=1}^{N}(|x_{i}|^{2}+\eta)^{\alpha/2}$. This leads to the following bound on the residual:
$\|\mathcal{A}\hat{\mathbf{x}}^{\dagger}-\mathbf{v}^{\dagger}\|_{2}\leq\sqrt{2\,\mathcal{L}(\hat{\mathbf{x}}^{\dagger},\gamma)}\leq\sqrt{2(\|\hat{\mathbf{x}}^{0}\|_{\alpha}^{2}+N)}.$ (84)
Now consider the recovery error with respect to a ground-truth $s$-sparse signal $\hat{\mathbf{x}}^{\diamondsuit}$ with support set $S:=\mathrm{supp}(\hat{\mathbf{x}}^{\diamondsuit})$. Let $S^{\dagger}$ be the index set corresponding to the $s$ largest entries (in magnitude) of $\hat{\mathbf{x}}^{\dagger}$. Using standard sparse recovery arguments and the RIP assumption $\delta_{2s}<1$, we have:
$\|\hat{\mathbf{x}}^{\dagger}-\hat{\mathbf{x}}^{\diamondsuit}\|_{2}\leq\|(\hat{\mathbf{x}}^{\dagger}-\hat{\mathbf{x}}^{\diamondsuit})_{S\cup S^{\dagger}}\|_{2}+\|(\hat{\mathbf{x}}^{\dagger})_{(S\cup S^{\dagger})^{c}}\|_{2}$
$\leq\frac{1}{\sqrt{1-\delta_{2s}}}\|\mathcal{A}(\hat{\mathbf{x}}^{\dagger}-\hat{\mathbf{x}}^{\diamondsuit})\|_{2}+\frac{1}{\sqrt{1-\delta_{2s}}}\|(\hat{\mathbf{x}}^{\dagger})_{(S\cup S^{\dagger})^{c}}\|_{2}$